# Leveraging Diverse Data Generation for Adaptable Zero-Shot Dialogue State Tracking

### v1

@@ Abstract
This work demonstrates that substantial gains in zero-shot dialogue state tracking (DST) accuracy can be achieved by increasing the diversity of training data using synthetic data generation techniques. Current DST training resources are severely limited in the number of application domains and slot types they cover due to the high costs of data collection, resulting in limited adaptability to new domains. The presented work overcomes this challenge using a novel, fully automatic data generation approach to create synthetic zero-shot DST training resources. Unlike previous approaches for generating DST data, the presented approach generates entirely new application domains to generate dialogues, complete with silver dialogue state annotations and slot descriptions. This approach is used to create the D0T dataset for training zero-shot DST models, which covers an unprecedented $1,000+$ domains. Experiments performed on the MultiWOZ benchmark indicate that training models on diverse synthetic data yields a performance improvement of $+6.7\%$ Joint Goal Accuracy, achieving results competitive with much larger models.
Leveraging Diverse Data Generation for
@@
James D. Finch  and Boxin Zhao  and Jinho D. Choi Department of Computer Science Emory University Atlanta, GA, USA {jdfinch, zinc.zhao, jinho.choi}@emory.edu
@@ 1 Introduction
A critical task for building task-oriented dialogue (TOD) systems is Dialogue State Tracking (DST), which aims to maintain a structured representation of the key task-related information provided throughout a dialogue. Conventionally, the state representation is composed of a set of task-specific slot-value pairs, where slots are information types provided by a predefined slot schema. While DST has been studied in fully supervised Heck et al. (2020); Xie et al. (2022); Won et al. (2023) and few-shot settings Lin et al. (2021); Shin et al. (2022); Chen et al. (2023), these settings rely on a substantial amount of labeled training examples within the targeted task domain. Recently zero-shot DST has gained more attention, since it requires the DST model to adapt to an unseen target domain for which no training examples are available Gupta et al. (2022); Wang et al. (2023); Heck et al. (2023).
Recently, it has been shown that an effective approach for zero-shot DST is to leverage slot descriptions to perform cross-task transfer Lin et al. (2021); Gupta et al. (2022); Zhao et al. (2022); Tavares et al. (2023). In this approach, a model is trained to interpret the slot descriptions to perform DST using gold supervision in several data-rich domains. During inference, the model can then interpret new slot descriptions to perform DST in unseen target domains without any training data. However, for this approach to succeed, sufficiently diverse training data must be available to enable the model to generalize and handle new slot types. We hypothesize that existing training data for DST is a bottleneck, as the two most popular datasets for DST training, MultiWOZ Budzianowski et al. (2018) and SGD Rastogi et al. (2020), only cover 5 and 16 domains respectively.
This work aims to explore the impact of increasing training data diversity on zero-shot DST performance. Since traditional methods of creating DST training data are costly and difficult to scale, we develop a novel, fully automatic data generation approach for zero-shot DST. This approach leverages the capabilities of instruction-tuned LLMs to create new task domains from scratch. Synthetic dialogues are generated for each domain, and are automatically annotated for dialogue state, complete with descriptions of labeled slots. This approach is leveraged to generate a synthetic DST dataset of unprecedented diversity, including over $1,000$ task domains. Experiment results demonstrate a substantial performance boost provided by this synthetic data on standard benchmarks. We make all models, code, and data publicly available to support future work. [^1] https://github.com/jdfinch/dot
In summary, our contributions are:
@@
    A novel approach for generating domain-diverse DST data
@@
    A synthetic DST dataset with $1,000+$ domains for training zero-shot models
@@
    Experimental results that demonstrate the impact of increasing training domain diversity on standard zero-shot DST benchmarks
@@ 2 Related Work
Current state-of-the-art (SoTA) approaches to zero-shot DST use sequence-to-sequence (S2S) formulations to predict appropriate values given a natural language specification of each slot to track Gupta et al. (2022); King and Flanigan (2023). This S2S modeling formulation has been shown to be a simple but effective tool for adapting to new slot types, since models can leverage descriptions of a new, unseen slot type via in-context learning (ICL) when making predictions. Recently, works using large language models (LLMs) have achieved state-of-the-art performance on zero-shot DST due to the excellent zero-shot ability of LLMs Hu et al. (2022); King and Flanigan (2023). However, since the cost of LLM decoding is too steep for many TOD applications, there are ongoing efforts to achieve competitive results with smaller models using cross-task transfer, where the model is trained on an existing set of task domains before being transferred to the unseen target domain Wang et al. (2023); Aksu et al. (2023).
@@
Successfully training a low-cost zero-shot DST model that generalizes to unseen domains depends on the quality and diversity of its training data; however, collecting a training resource that covers many TOD domains is costly. The most popular dataset, MultiWOZ, was collected using a wizard-of-oz setup using human participants and only covers 5 domains Budzianowski et al. (2018). The Schema Guided Dialogues (SGD) dataset was created in an attempt to increase the diversity of available DST resources using a rule-based data generation approach, where the final dialogue text was paraphrased by crowdworkers to improve naturalness Rastogi et al. (2020). Even with this more cost-effective collection technique, SGD only covers 16 domains in its training split. Moreover, both datasets suffer from high inter-domain similarity. In the case of MultiWOZ, each domain covers a component of a travel planning application, in which a user talks to an artificial travel agent. As a result, there is a high degree of topical and structural similarity between dialogues, and all domains share a similar focus on scheduling. This results in many overlapping slots between domains to cover scheduling details such as dates, times, and locations. SGD has a more diverse array of domains, yet most are similar to MultiWOZ in that they focus on booking and scheduling. In particular, the Bus, Calendar, Event, Flight, Hotel, RentalCar, Service, and Train domains all share this scheduling focus. As a result of this limited diversity and the cost of additional data collection, it is unknown whether the domain coverage of existing DST resources is a bottleneck for training a zero-shot DST model with robust cross-task transfer.
@@
Several previous works explore data augmentation methods for improving the diversity of limited DST data. Nearly all of these approaches target the few-shot setting, where a limited [^2] Typical few-shot DST evaluation settings use at least several hundred labeled dialogue examples as ”few” shots. number of labeled examples are used as a seed set to be augmented with additional, synthetic examples. This can be done using simple approaches to improve the lexical Quan and Xiong (2019); Yin et al. (2020) or semantic Summerville et al. (2020); Lai et al. (2022) diversity of training examples, or by synthesizing entire dialogues Campagna et al. (2020); Aksu et al. (2021, 2022); Mehri et al. (2022); Mohapatra et al. (2021); Kim et al. (2021); Wan et al. (2022) to create additional training resources. These previous works in DST data generation demonstrate that automatic methods for data augmentation and generation can help address the limitations of existing training resources and improve transfer to data-poor domains. Additional detail regarding related work in DST data generation is provided in Appx. A.
Our presented data generation approach is distinct from all previous DST data generation methods because it generates entirely new task domains, in addition to new dialogues with silver annotations. Furthermore, our approach is fully automatic, requiring no few-shot data or manual creation of domain-specific resources, making it ideal for scaling up the diversity of training resources for zero-shot DST.
3 Generating DST Data
This section presents our fully automatic data generation approach to support training DST models capable of zero-shot domain transfer. The goal of the approach is to create a set of dialogue data covering many diverse task domains, with silver dialogue state labels and natural language slot descriptions. Given the exceptional zero-shot performance of instruction-tuned large language models (LLMs) on a wide variety of tasks Brown et al. (2020); Kojima et al. (2022); Heck et al. (2023), our approach explores using instruction-tuned LLMs for data generation. We use GPT333gpt-3.5-turbo-0301 is used for all stages of the approach, except for QA Pair Generation in which gpt-4-0314 is used. in all of our presented experiments, although any LLM can be used for our approach in principle.
The approach consists of four stages, which are summarized in Figure 1. First, domains are derived through an iterative process of generating and refining dialogue scenario descriptions (§3.1). Next, a dialogue is crafted based on the scenario description and a generated unstructured information list corresponding to the scenario (§3.2). Third, each turn in each dialogue is automatically annotated with silver dialogue state labels (§3.3). Finally, a slot description is composed for each silver slot-value pair annotation (§3.4). All prompts included in the approach are provided in Appx. B.
@@ 4       $\mathbb{E}\leftarrow\textbf{cluster}(\textbf{embed}(S\cup S^{\prime}))$
Algorithm 1 shows our scenario derivation method. GPT is iteratively prompted to create a mini-set of $k$ dialogue scenario descriptions (L3). Each mini-set is combined with the scenarios obtained from previous iterations, where each scenario description is encoded into an embedding by SentenceBERT [^4] SentenceBert model: all-MiniLM-L6-v2 Reimers and Gurevych (2019) and the resulting embeddings are clustered through a community detection algorithm (L4). [^5] https://www.sbert.net/docs/package_reference/util.html A deduplicated set of scenario descriptions is created by selecting one embedding from every cluster, which is mapped back to its corresponding scenario description (L5). This iteration continues until the set reaches the requested size (L2). In our case, $k=100,n=1000$. A sample of the generated scenarios is given in Appx. C.
@@ 3.4 Slot Description Generation
For each state update $U_{t}$ produced by automatic annotation (§3.3), GPT is instructed to generate a specification of each slot in $U_{t}$ using a single prompt. The prompt includes each slot value pair $(s^{t}_{i},v^{t}_{i})$ in $U_{t}$ as well as each question $q^{t}_{i}$ corresponding to each slot. GPT is asked to generate a specification for each slot as a short natural language description $d^{t}_{i}$, in addition to a few comma-separated example values $e^{t}_{i}$ that could fill the slot.
@@ 4 New Dataset for Zero-Shot Tracking
Slot-Values 324,973 TokensSV 2.0
@@
Using our DST data generation approach (§3), we create a Diverse 0-shot Tracking dataset: D0T. Since we aim to measure the impact of increasing the diversity of DST training resources, we generate D0T to include an unprecedented $1,000+$ domains and $5$ dialogues per domain. Applying automatic state annotation (§3.3) to the generated dialogues yields $324,973$ slot-value pairs in state updates. Since compiling each dialogue state $S_{t}=update(S_{t-1},U_{t})$ produces an excessive $\approx 6.5$ million total slot-value pairs for DST training, slot-value pairs are downsampled using a method that maintains slot type diversity. We randomly sample exactly $1$ example for each of the original $324,973$ slot-value updates from the set of final slot-values where that slot is filled (non-empty), resulting in $n=324,973$ filled slot-value examples. To include examples of empty slots, we randomly sample $m$ empty slot-value pairs from the final compiled states, where $m=0.5*n$. Table 1 presents the final statistics of the dataset. We validate the quality of the dataset by recruiting 3 human evaluators to annotate $60$ randomly sampled turns, judging (1) whether each slot-value pair correctly represents information in the corresponding turn and (2) whether each state update $U_{t}$ is missing any important information in the turn. $82.0\%$ of slot-value pairs were judged correct and only $7\%$ of state updates were missing important information.
@@ 5 Experiment Setup
Zero-Shot DST experiments follow the standard setup, using the MultiWOZ data in a leave-one-out evaluation design Budzianowski et al. (2018). MultiWOZ 2.4 Ye et al. (2022) is used as it contains corrected gold labels in the validation and test splits, and an evaluation on MultiWOZ 2.1 Eric et al. (2020) is also included to facilitate further comparison to previous work.
@@
Two base models, T5-3B Raffel et al. (2020) and Llama2-13B-Chat [^6] Llama2 training uses QLoRA Dettmers et al. (2023). Ouyang et al. (2022), are trained using the sequence-to-sequence format shown in Figure 3, which uses the "independent" formulation from Gupta et al. (2022).
@@
A: Good afternoon, Mr. Smith. I’m here today to survey your land and assess its value. B: Of course, please go ahead. A: Firstly, can you tell me the location and size of the land? B: Sure. The land is located on the outskirts of town, about 10 miles away from the city center. It’s approximately 20 acres. A: That’s helpful. Can you also tell me about the type of terrain and land features on the property? Identify the information from the above dialogue: land size: the area encompassed by the property, typically measured in units such as acres, hectares, or square miles. (e.g. 50 hectares, 2 square miles)?
@@
Figure 3: An example of an input token sequence from the D0T dataset used for training. [YELLOW]: dialogue context $D_{1..t}$   [PEACH]: slot $s^{t}_{i}$   [GREEN]: slot description $d^{t}_{i}$   [RED]: value examples $e^{t}_{i}$
Models are trained in two stages. First, the model is trained on the D0T dataset in order to optimize domain adaptability. Since D0T is a synthetic dataset with some noise in its annotations, the model is then refined using the standard training split of the MultiWOZ data (excluding the target domain being evaluated). Baseline models are trained on the MultiWOZ training only, identical to previous work on zero-shot DST.

### v2

# Diverse and Effective Synthetic Data Generation for Adaptable Zero-Shot Dialogue State Tracking

@@ Abstract
We demonstrate substantial performance gains in zero-shot dialogue state tracking (DST) by enhancing training data diversity through synthetic data generation. Existing DST datasets are severely limited in the number of application domains and slot types they cover due to the high costs of data collection, restricting their adaptability to new domains. This work addresses this challenge with a novel, fully automatic data generation approach that creates synthetic zero-shot DST datasets. Distinguished from previous methods, our approach can generate dialogues across a massive range of application domains, complete with silver-standard dialogue state annotations and slot descriptions. This technique is used to create the D0T dataset for training zero-shot DST models, encompassing an unprecedented 1,000+ domains. Experiments on the MultiWOZ benchmark show that training models on diverse synthetic data improves Joint Goal Accuracy by 6.7%, achieving results competitive with models 13.5 times larger than ours.
Diverse and Effective Synthetic Data Generation for
@@
James D. Finch  and Jinho D. Choi Department of Computer Science Emory University Atlanta, GA, USA {jdfinch, jinho.choi}@emory.edu
@@ 1 Introduction
A critical task for building task-oriented dialogue (TOD) systems is Dialogue State Tracking (DST), which aims to maintain a structured representation of the key task-related information provided throughout a dialogue. Conventionally, the state representation is composed of a set of task-specific slot-value pairs, where slots are information types provided by a predefined slot schema. While DST has been studied in fully supervised Heck et al. (2020); Xie et al. (2022); Won et al. (2023) and few-shot settings Lin et al. (2021); Shin et al. (2022); Chen et al. (2023), these settings rely on a substantial amount of labeled training examples within the targeted task domain. To this end, zero-shot DST has recently gained attention, as it requires the DST model to adapt to an unseen target domain for which no training examples are available Gupta et al. (2022); Wang et al. (2023); Heck et al. (2023).
Leveraging slot descriptions to perform cross-task transfer is shown to be effective for zero-shot DST Lin et al. (2021); Gupta et al. (2022); Zhao et al. (2022); Tavares et al. (2023). In this approach, a model is trained to interpret the slot descriptions to perform DST using gold supervision in several data-rich domains. During inference, the model interprets new slot descriptions to perform DST in unseen target domains without any training data. However, for this approach to succeed, sufficiently diverse training data must be available to enable the model to generalize and handle new slot types. We hypothesize that existing training data for DST is a bottleneck, as the two most popular datasets for DST training, MultiWOZ Budzianowski et al. (2018) and SGD Rastogi et al. (2020), only cover 7 and 16 domains, respectively.
This work aims to explore the impact of increasing training data diversity on zero-shot DST performance. Since traditional methods of creating diverse DST training data are costly and difficult to scale, we develop a novel, fully automatic data generation approach for zero-shot DST. This approach leverages the capabilities of instruction-tuned large language models (LLMs) to create new task domains from scratch. Synthetic dialogues are generated for each domain, and are automatically annotated for dialogue state, complete with descriptions of labeled slots. This approach is leveraged to generate a synthetic DST dataset of unprecedented diversity, including over $1,000$ task domains. Experiment results demonstrate a substantial performance boost provided by this synthetic data on standard benchmarks. In summary, our contributions are:
@@
    A novel approach for generating domain-diverse DST data.
@@
    A synthetic DST dataset with $1,000+$ domains for training zero-shot models.
@@
    Efficient state-of-the-art models that robustly handle diverse domains for zero-shot DST.
We make all models, code, and data publicly available to support future work. [^1] https://github.com/anonymous
@@ 2 Related Work
Current state-of-the-art (SoTA) approaches to zero-shot DST use sequence-to-sequence (S2S) modeling to predict appropriate values given a natural language specification of each slot to track Gupta et al. (2022); King and Flanigan (2023). Such S2S modeling has been effective for adapting to new slot types, since models can leverage descriptions of a new, unseen slot type via in-context learning (ICL) when making predictions. Recently, models using LLMs have achieved state-of-the-art results on this task due to the excellent zero-shot ability of LLMs Hu et al. (2022b); King and Flanigan (2023). However, the cost of LLM decoding is often too steep for many task-oriented dialogue (TOD) applications. Thus, ongoing work aims to achieve SoTA results with smaller models using cross-task transfer, where the model is trained on an existing set of task domains before being transferred to the unseen target domain Wang et al. (2023); Aksu et al. (2023).
@@
Successful modeling of a low-cost zero-shot DST model that generalizes to unseen domains depends on the quality and diversity of its training data; however, collecting a training resource that covers diverse TOD domains is costly. The most popular dataset, MultiWOZ, was collected using a wizard-of-oz setup using human participants, yet only covers 7 domains Budzianowski et al. (2018). The Schema Guided Dialogues (SGD) dataset was created in an attempt to increase the diversity of available DST resources using a rule-based data generation approach, where the final dialogue text was paraphrased by crowdworkers to improve naturalness Rastogi et al. (2020). Even with this more cost-effective collection technique, SGD only covers 16 domains in its training split. Moreover, both datasets suffer from high inter-domain similarity. In the case of MultiWOZ, each domain covers a component of a travel planning application, in which a user talks to an artificial travel agent. As a result, there is a high degree of topical and structural similarity between dialogues, and all domains share a similar focus on scheduling. This results in many overlapping slots between domains to cover scheduling details such as dates, times, and locations. SGD has a more diverse array of domains, yet most are similar to MultiWOZ in that they focus on booking and scheduling. In particular, the Bus, Calendar, Event, Flight, Hotel, RentalCar, Service, and Train domains all share this scheduling focus. As a result of this limited diversity and the cost of additional data collection, it is unknown whether the domain coverage of existing DST resources is a bottleneck for training a zero-shot DST model with robust cross-task transfer.
@@
Several previous works explore data augmentation methods for improving the diversity of limited DST data. Nearly all of these approaches target the few-shot setting, where a limited number of labeled examples are used as a seed set to be augmented with additional, synthetic examples. This can be done using simple approaches to improve the lexical Quan and Xiong (2019); Yin et al. (2020) or semantic Summerville et al. (2020); Lai et al. (2022) diversity of training examples, or by synthesizing entire dialogues Campagna et al. (2020); Aksu et al. (2021, 2022); Mehri et al. (2022); Mohapatra et al. (2021); Kim et al. (2021); Wan et al. (2022) to create additional training resources. These previous works in DST data generation demonstrate that automatic methods for data augmentation and generation can help address the limitations of existing training resources and improve transfer to data-poor domains. Additional detail regarding related work in DST data generation is provided in Appendix A.
Our DST data generation approach is distinct from all previous methods because it generates entirely new task domains, in addition to new dialogues with silver annotations. Furthermore, our approach is fully automatic, requiring no few-shot data or manual creation of domain-specific resources, making it ideal for scaling up the diversity of training resources for zero-shot DST.
3 DST Data Generation
This section presents our fully automatic data generation approach to support training DST models capable of zero-shot domain transfer. Our goal is to create a set of dialogue data covering many diverse task domains, with silver dialogue state labels and natural language slot descriptions. Given the exceptional zero-shot performance of instruction-tuned large language models (LLMs) on a wide variety of tasks Brown et al. (2020); Kojima et al. (2022); Heck et al. (2023), our approach explores using instruction-tuned LLMs for data generation. We use GPT222gpt-3.5-turbo-0301 is used for all stages of the approach, except for QA Pair Generation in which gpt-4-0314 is used. in all of our presented experiments, although any LLM can be used for our approach in principle.
The approach consists of four stages, which are summarized in Figure 1. First, domains are derived through an iterative process of generating and refining dialogue scenario descriptions (§3.1). Next, a dialogue is crafted based on the scenario description and a generated unstructured information list corresponding to the scenario (§3.2). Third, each turn in each dialogue is automatically annotated with silver dialogue state labels (§3.3). Finally, a slot description is composed for each silver slot-value pair annotation (§3.4). All prompts included in the approach are provided in Appendix B.
@@ 3.1 Scenario Derivation
Algorithm 1 shows our scenario derivation method. GPT is iteratively prompted to create a mini-set of $k$ dialogue scenario descriptions (L3). Each mini-set is combined with the scenarios obtained from previous iterations, where each scenario description is encoded into an embedding by SentenceBERT [^3] SentenceBert model: all-MiniLM-L6-v2 Reimers and Gurevych (2019) and the resulting embeddings are clustered through a community detection algorithm (L4). [^4] https://www.sbert.net/docs/package_reference/util.html A deduplicated set of scenario descriptions is created by selecting one embedding from every cluster, which is mapped back to its corresponding scenario description (L5). This iteration continues until the set reaches the requested size (L2). In our case, $k=100,n=1000$. Appx. C gives a sample of the generated scenarios.
@@ 3.4 Slot Description Generation
For each state update $U_{t}$ produced by automatic annotation (§3.3), GPT is instructed to generate a specification of each slot in $U_{t}$ using a single prompt. The prompt includes each slot value pair $(s^{t}_{i},v^{t}_{i})$ in $U_{t}$ as well as each question $q^{t}_{i}$ corresponding to each slot. GPT is asked to generate a description for each slot as a short natural language phrase $d^{t}_{i}$, in addition to a few comma-separated example values $e^{t}_{i}$ that could fill the slot.
@@ 4 New Dataset for Zero-Shot Tracking
Using our DST data generation approach (§3), we create a Diverse 0-shot Tracking dataset: D0T. Since we aim to measure the impact of increasing the diversity of DST training resources, we generate D0T to include unprecedented $1,000+$ domains and $5$ dialogues per domain. Applying automatic state annotation (§3.3) to the generated dialogues yields $324,973$ slot-value pairs in state updates. Since compiling each dialogue state $S_{t}=update(S_{t-1},U_{t})$ produces an excessive $\approx 6.5$ million total slot-value pairs for DST training, slot-value pairs are downsampled using a method that maintains slot type diversity. We randomly sample exactly $1$ example for each of the original $324,973$ slot-value updates from the set of final slot-values where that slot is filled (non-empty), resulting in $n=324,973$ filled slot-value examples. To include examples of empty slots, we randomly sample $m$ empty slot-value pairs from the final compiled states, where $m=0.5*n=162,487$. Table 1 presents the final statistics of the dataset, and Table 2 presents a comparison to existing data.
@@
Slot-Values 487,460 TokensSV 2.0
@@
We validate the quality of the dataset by recruiting 3 human evaluators to annotate $60$ randomly sampled turns, judging (1) whether each slot-value correctly represents information in the corresponding turn and (2) whether each state update $U_{t}$ is missing any important information in the turn. $82\%$ of slot-value pairs were judged correct and $7\%$ of state updates were missing important information.
Dataset Dom. Dial. Turns SV US
MWOZ 7 8,438 113,556 4,510 24
SGD 16 16,142 329,964 14,139 214
D0T 1,003 5,015 100,471 487,460 173,572
Table 2: Comparison of D0T to the train splits of MultiWOZ 2.1/2.4 (MWOZ) and SGD, compared on number of domains (Dom.), dialogues (Dial.), turns, slot-values (SV), and unique slot names (US).
@@ 5 Experiment Setup
Evaluation Data
Our experiments on zero-shot DST use the standard MultiWOZ benchmark Budzianowski et al. (2018). This evaluation was designed using a leave-one-out setup in which a zero-shot DST model is tested on each of five domains (Attraction, Hotel, Restaurant, Taxi, Train) after being trained on the other four, to test zero-shot transfer to new domains. Joint Goal Accuracy (JGA) is the evaluation metric, measuring the proportion of turns for which the entire dialogue state is correctly inferred. The MultiWOZ 2.4 Ye et al. (2022) variant is used as the main evaluation dataset since it contains corrected gold labels in the validation and test splits. We additionally include an evaluation on the uncorrected MultiWOZ 2.1 variant Eric et al. (2020) to facilitate further comparison to previous work.
Since MultiWOZ does not contain slot descriptions, a single-sentence description is written for each MultiWOZ slot to provide slot definitions. Descriptions are authored based on Lin et al. (2021) but with improvements in detail and grammar. Additionally, descriptions are augmented with $4$ value examples for each slot. No prompt engineering or validation experiments are performed when creating slot descriptions and value examples, to reflect the performance of the model in real-world settings without requiring extensive development effort.
Models
The impact of domain-diverse training data on zero-shot DST is evaluated by comparing models that leverage the domain-diverse D0T dataset as a training resource against baselines trained only on the standard training splits of benchmark data. Models leveraging D0T (+D0T) are trained in two sequential training stages. Models are first trained on D0T to acquire domain-general state tracking ability, and then refined in a second training stage using the standard training split of benchmark data.
@@
Two base models, T5 1.1 Raffel et al. (2020) and Llama2-Chat Ouyang et al. (2022), are used in our experiments. We use the 11B and 13B variants of the T5 and Llama2 models, respectively; however, for greater efficiency and robustness for two-stage model training, we additionally leverage the QLoRA Dettmers et al. (2023) quantization and training method. Models are trained using the sequence-to-sequence format shown in Figure 3 which follows the "independent" formulation from Gupta et al. (2022). Appendix D provides implementation details such as model hyperparameters.
@@
A: Good afternoon, Mr. Smith. I’m here today to survey your land and assess its value. B: Of course, please go ahead. A: Firstly, can you tell me the location and size of the land? B: Sure. The land is located on the outskirts of town, about 10 miles away from the city center. It’s approximately 20 acres. A: That’s helpful. Can you also tell me about the type of terrain and land features on the property? Identify the information from the above dialogue: land size: the area encompassed by the property, typically measured in units such as acres, hectares, or square miles. (e.g. 50 hectares, 2 square miles)? ex. The floodwaters have submerged over 150 hectares      of farmland. land size? -> 150 hectares ex. Yes, we’re finalizing a purchase of 50 acres in the valley.      land size? -> 50 acres
@@
Figure 3: An example of an input token sequence from the D0T dataset used for training. [YELLOW]: dialogue context $D_{1..t}$   [PEACH]: slot $s^{t}_{i}$   [GREEN]: slot description $d^{t}_{i}$   [RED]: value examples $e^{t}_{i}$   [BLUE]: In-context demonstrations (+ICL only)
Additionally, since recent work in zero-shot DST has shown performance improvements from including demonstrations in slot descriptions using in-context learning Gupta et al. (2022); Hu et al. (2022b); King and Flanigan (2023), we also experiment with this approach using the Llama2 base model, to observe the interaction between domain-diverse training and in-context demonstration. Models leveraging in-context demonstrations (+ICL) are trained and tested with slot descriptions that include up to $k=3$ in-context demonstrations, where $k$ is a per-domain hyperparameter selected by validation performance.
For MultiWOZ, demonstrations are collected for each slot by manually constructing $3$ single-turn examples of the slot being updated with an appropriate value. For D0T, we collect in-context demonstrations using a fully automatic method in order to preserve the fully-automatic nature of the data generation approach. This is done by augmenting slot descriptions in the D0T dataset by sampling slot-value labels that share similar semantics to the target slot. Similar slot-value examples are found for demonstration sampling by encoding every silver slot-value update label in D0T as the token sequence "$s\text{: }v$" using SBERT Reimers and Gurevych (2019) and then clustering the encoded slot-values using HDBSCAN McInnes et al. (2017). Then, for each training example of slot name, value, and slot description $(s,v,d)$, up to $3$ demonstrations are randomly sampled from other training examples that appear in the same cluster and the same domain, but different dialogues. The description $d$ is augmented by appending each sampled demonstration value with the text of the dialogue turn in which it appears, using the format exemplified in Figure 3.
6 Results